OpenStudy (anonymous):

I am having the most difficult time trying to understand a problem with the formula a=pe^rt, where A=amount accumulated, p=principal, r=interest rate, and t=time in yrs. What principal should be deposited at 7.405% to ensure the account will be worth $18,000 in 11 years

OpenStudy (anonymous):

18,000 = p(1+.07405)^11

OpenStudy (anonymous):

So 18,000=(1p+7405p)^11, right?

OpenStudy (anonymous):

First step would be to add (1 + .07405) to get 1.07405 Then take that to the ^11

OpenStudy (anonymous):

2.194166065

OpenStudy (anonymous):

I thought the formula was A= p/e^rt

OpenStudy (anonymous):

Good now you have 18,000 = 2.194166065p so divide both sides by 2.194166065

OpenStudy (anonymous):

8203.572322= p

OpenStudy (anonymous):

or $8203.57?

OpenStudy (anonymous):

OK.. we can do it your way... just a minute.

OpenStudy (anonymous):

18,000 = P e^(.07405*11) 18,000 = Pe^.81455 18,000 = P 2.258 Divide by 2.258 $7971.66

OpenStudy (anonymous):

How do you calculate e?

OpenStudy (anonymous):

I plugged it into the calculator... I have an e^x key.

OpenStudy (anonymous):

Wait just a minute

OpenStudy (anonymous):

Also the ln key right?

OpenStudy (anonymous):

Ok, I got the same thing you got this time.

OpenStudy (anonymous):

I am not used to the e function.. here is an example, I will do yours for you following this example So 130000 = 1949 e^(.00835 t) Take logs of both sides: ln 130000 = ln 1949 e^(.00835 t) = ln 1949 + ln e^(.00835 t) = ln 1949 + .00835 t t = (ln 130000 - ln 1949) / .00835 = (ln 130000/1949) / .00835 which you have to calculate using a calculator

OpenStudy (anonymous):

18,000 = Pe^(.07405 * 11) 18,000 = Pe^.81455 ln 18000 = ln P + ln(e^.81455) ln 18000 = ln P + .81455 9.79813 - .81455 = ln P 8.98358 = ln P e^8.98358 = e^ (ln P) $7971.12 = P

OpenStudy (anonymous):

ok. It was the "e" part that was throwing me off. But what you and I got is what the book has also.

OpenStudy (anonymous):

Great... A little out of my league... but I am glad we got it.

OpenStudy (anonymous):

Now, with the formula with A =p(1+r/n)^nt A = $982500 p+71500, t=17yr

OpenStudy (anonymous):

Don't we have to know either r or n?

OpenStudy (anonymous):

I have 98, 2500=71, 500(1+r/n)^17. Do you know how I would find r/n?

OpenStudy (anonymous):

R is what they are asking for. R is a percent

OpenStudy (anonymous):

What is n?

OpenStudy (anonymous):

Here's the question: For compound interest accts, the amount A accumulated or due depends on the principal p, interest rate r, number of compounding per year n, and the time t in years according to the formula A=p(1 + r/n)^nt. Find the r given A=$98, 250, p=71,500, and t=17 yr with interest compound monthly

OpenStudy (anonymous):

compounded monthly means that n = 12

OpenStudy (anonymous):

982,500 = 71500(1 + r/12)^12*17

OpenStudy (anonymous):

oh, because 12 represent the "per year"

OpenStudy (anonymous):

982,500 = 71,500(1 + r/12)^204

OpenStudy (anonymous):

divide both sides by 71,500

OpenStudy (anonymous):

13.7412587 = (1 + r/12)^204

OpenStudy (anonymous):

13. 74128874=(1+r/12)^204

OpenStudy (anonymous):

Now you have to get rid of the ^204 you do that by taking the 204th root of both sides. Use the x sqrt(y) key on your calculator

OpenStudy (anonymous):

1.012927965 = 1 + r/12

OpenStudy (anonymous):

Subtract 1 from both sides

OpenStudy (anonymous):

.012927965 = r/12

OpenStudy (anonymous):

multiply both sides by 12

OpenStudy (anonymous):

12.155 = r

OpenStudy (anonymous):

ON the step with the sq rt of 13.7412587 I got an error msg

OpenStudy (anonymous):

Your saying 204 sq rt 13.7412587 correct?

OpenStudy (anonymous):

You are not taking the sqrt root.. Find the key that has x sqrt(y) type in 13.7412587 hit the xsqrt(y) key and then type in 204 and hit enter

OpenStudy (anonymous):

I dont believe I have that key

OpenStudy (anonymous):

Are you operating a TI-84?

OpenStudy (anonymous):

NO.. just a calculator on my computer the basic one..

OpenStudy (anonymous):

Let me see if I can find a TI-84 on line and tell you what it looks like

OpenStudy (anonymous):

How about under MATH option 5 does it show an x and then a radical sign

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You want the x to be the 204 and under the radical sign to be the 13.xxxxxx Try it and see if you can get what I got.

OpenStudy (anonymous):

YES!!!! I got it! such a relief! thank you very much